OpenStudy (anonymous):

Verify this trigonometric equation help please! -tan^2x+sec^2x=1

OpenStudy (anonymous):

well you can do this..... tan^2x = (sin^2x/cos^2x) and remember that sec^2x = 1/cos^2x. just sub and do some algebra.. give it a try @jtschellenberg

OpenStudy (anonymous):

ok i did and now i'm at (sin^2x+1)/cos^2x what next?

OpenStudy (anonymous):

@mebs

OpenStudy (callisto):

It should be (-sin^2x+1)/cos^2x instead of (sin^2x+1)/cos^2x

OpenStudy (anonymous):

wait no can someone explain?

OpenStudy (anonymous):

so here it is -sin^2x/cos^2x + 1/cos^2x = 1

OpenStudy (anonymous):

it has common denominators so -sin^2x +1/cos^2x =1

OpenStudy (anonymous):

right. what next?

OpenStudy (callisto):

\[LS\]\[=-tan^2x+sec^2x\]\[=-\frac{sin^2x}{cos^2x}+\frac{1}{cos^2x}\]\[=\frac{-sin^2x+1}{cos^2x}\]For the numerator, use the identity \(sin^2x+cos^2x=1\)

OpenStudy (anonymous):

now recall that sin^2x +cos^2x = 1 right

OpenStudy (anonymous):

yeah i remember that one

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/4112/verify-this-trigonometric-equation-help-please ask your trigonometric homework questions at http://www.mathskey.com/question2answer/ … and get free math help. all the best

OpenStudy (anonymous):

just substitute -sin^2x +1 by cos^2x and wala they cancel and you get 1

OpenStudy (anonymous):

oh so -sin^2x+1=cos^2x?

OpenStudy (anonymous):

yes